# Beyond Blind Spots: Analytic Hints for Mitigating LLM-Based Evaluation Pitfalls

### v1

@@ Abstract
Experiments on a test set of 100 programs using four production-level LaaJs show that LaaJ alone detects only about 45% of the errors present in the code (in all judges we tested), while the analytic checker alone lacks explanatory depth. When combined, the LaaJ+Hints configuration achieves up to 94% coverage (for the best-performing judge and injection prompt) and produces qualitatively richer, more accurate explanations, demonstrating that analytic–LLM hybrids can substantially enhance evaluation reliability in deployed pipelines. We release the dataset and all used prompts.
@@ Introduction
LLMs are increasingly used as automated judges in code generation (zhuo-2024-ice; tong-zhang-2024-codejudge; Zhao2024CodeJudgeEvalCL). However, recent studies have shown that LaaJs often overlook fine-grained errors and struggle to capture semantic correctness, particularly in complex domains like code generation (doddapaneni2024-fbi; tong-zhang-2024-codejudge; tan2024-judgebench). To address these issues, several works proposed enhancing LaaJ reasoning with prompt-level interventions, such as self-refinement (MadaanSelfRefine), rationale rewriting (trivedi2024self), or tool-assisted evaluation (toolformer).
Closest to our approach are hint-injection methods that steer LLM reasoning with targeted cues, including AutoHint (learned task-specific hints), Directional Stimulus Prompting (policy-generated stimulus tokens) (stimulusPrompting), Progressive-Hint Prompting (reusing prior answers as hints) (zheng2023progressive), and Hint-before-Solving (pre-drafted hints) (Fu2024HintbeforeSolvingPG).
@@ Results
As shown in Figure 3, introducing analytic hints led to consistent gains across all models. Native judges detected roughly 45–53 % of total issues, whereas hybrid judges reached 63–94.4%, depending on model and prompt. Most of the improvement stems from capturing analytic-only errors previously missed by the native judges, while performance on native-only and overlap categories remained stable or improved slightly. The optimized hint injection achieved the highest coverage for all but one model, with DeepSeek-v3 and gpt-oss-120b exceeding 90 %. Overall, analytic hint injection yielded a 1.5–2× increase in total detection coverage.
@@ Conclusions
We presented a practical approach for enhancing LaaJs using analytic hint injection in code evaluation. Grounded in expert analysis of COBOL evaluation failures, our method couples a taxonomy of blind spots with a lightweight checker that emits targeted hints. Injecting these hints into the judge’s prompt refocuses its reasoning toward previously overlooked issues, yielding significant gains without retraining. Among all models, the gpt-oss-120b judge with hints achieved the best performance, addressing 94.4% of errors while retaining 67.31% of those identified by the native judge. This demonstrates that prompt-level analytic interventions can substantially improve judgment coverage while preserving general evaluative capabilities.

### v2

@@ Abstract
Experiments on a test set of 100 programs using four production-level LaaJs show that LaaJ alone detects only about 45-63% of the errors present in the code (in all judges we tested), while the analytic checker alone lacks explanatory depth. When combined, the LaaJ+Hints configuration achieves up to 74% coverage (for the best-performing judge and injection prompt) and produces qualitatively richer, more accurate explanations, demonstrating that analytic–LLM hybrids can substantially enhance evaluation reliability in deployed pipelines. We release the dataset and all used prompts.
@@ Introduction
LLMs are increasingly used as automated judges in code generation (Zhuo 2024; Tong and Zhang 2024; Zhao et al. 2024). However, recent studies have shown that LaaJs often overlook fine-grained errors and struggle to capture semantic correctness, particularly in complex domains like code generation (Doddapaneni et al. 2024; Tong and Zhang 2024; Tan et al. 2024). To address these issues, several works proposed enhancing LaaJ reasoning with prompt-level interventions, such as self-refinement (Madaan et al. 2023), rationale rewriting (Trivedi et al. 2024), or tool-assisted evaluation (Schick et al. 2023).
Closest to our approach are hint-injection methods that steer LLM reasoning with targeted cues, including AutoHint (learned task-specific hints), Directional Stimulus Prompting (policy-generated stimulus tokens) (Li et al. 2023), Progressive-Hint Prompting (reusing prior answers as hints) (Zheng et al. 2023), and Hint-before-Solving (pre-drafted hints) (Fu et al. 2024).
@@ Results
As shown in Figure 3, introducing analytic hints led to consistent gains across all models. Native judges detected roughly 45–63% of total issues, whereas hybrid judges reached 48–74.1%, depending on model and prompt. Most of the improvement stems from capturing analytic-only errors previously missed by the native judges, while performance on native-only and overlap categories remained stable or improved slightly. We observe that the optimized hint injection achieves consistently high success rates of issues detected with analytic checker and confirmed by LaaJ, with DeepSeek-v3 and gpt-oss-120b exceeding 90%.
@@ Conclusions
We presented a practical approach for enhancing LaaJs using analytic hint injection in code evaluation. Grounded in expert analysis of COBOL evaluation failures, our method couples a taxonomy of blind spots with a lightweight checker that emits targeted hints. Injecting these hints into the judge’s prompt refocuses its reasoning toward previously overlooked issues, yielding significant gains without retraining. Among all models, the gpt-oss-120b judge with hints achieved the best performance, addressing 74.1% of errors while retaining 67.31% of those identified by the native judge. This demonstrates that prompt-level analytic interventions can substantially improve judgment coverage while preserving general evaluative capabilities.